OpenStudy (anonymous):

integrate xsqrt(4-x) dx by using integration by parts method.

OpenStudy (anonymous):

\[\int x \times \sqrt{4-x} dx = x \int \sqrt{4-x} dx - \int [ \int \sqrt{4-x}dx \times \frac{dx}{dx}]\]

OpenStudy (anonymous):

\[x \times \frac{2}{3} (4 -x)^{\frac{3}{2}} - \frac{2}{3} \times \frac{2}{5} \times (4-x)^{\frac{5}{2}}\]

OpenStudy (anonymous):

the answers is \[2/3 (4-x)^{5/2} - 8/3 (4-x)^{3/2} +c\]

OpenStudy (anonymous):

okay I am back let me do this again

OpenStudy (anonymous):

ah you do remember the formula don't you

OpenStudy (anonymous):

yeah i do i tried the question with v = (4-x)^1/2 and u = x

OpenStudy (anonymous):

it gets me the wrong answer

OpenStudy (anonymous):

i used u substitution by letting u= 4-x and i get the right answer but its not using the method that i want to lol

OpenStudy (anonymous):

\[x \int \sqrt{4-x} dx - \int[ \int\sqrt{4-x}dx \times \frac{dx}{dx}]\]

OpenStudy (anonymous):

ill brb :)

OpenStudy (anonymous):

\[-x \times \frac{(4-x)^{1 + \frac{1}{2}}}{1 + \frac{1}{2}} - \int \frac{(4-x)^{1 + \frac{1}{2}}}{1 + \frac{1}{2}}dx\]

OpenStudy (anonymous):

\[-x \times \frac{(4-x)^{\frac{3}{2}}}{\frac{3}{2}} - \frac{1}{\frac{3}{2}} \times \frac{(4-x)^{\frac{3}{2} + 1 }}{ 1 + \frac{3}{2}} \]

OpenStudy (anonymous):

\[ -x \times \frac{2}{3} \times (4-x)^{\frac{3}{2}} - \frac{2}{3} \times \frac{2}{5} \times (4-x)^{\frac{5}{2}}\]

OpenStudy (anonymous):

ah I did that again I am still getting the same with a sign reverse

OpenStudy (anonymous):

ah Let's try this again with different approach

OpenStudy (anonymous):

\[4- x = t \] \[-dx = dt\]

OpenStudy (anonymous):

\[x = 4-t\]

OpenStudy (anonymous):

\[\int ( 4- t )(-1)(\sqrt t) dt\]

OpenStudy (anonymous):

\[\int (t - 4 ) \sqrt t dt \]

OpenStudy (anonymous):

\[ \int t^{1 + \frac{1}{2}} - 4 \times t^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[\int t^{1 + \frac{1}{2}} - 4 \times t^{\frac{1}{2}}dt\]

OpenStudy (anonymous):

\[\int t^{\frac{3}{2}}dt - 4 \int t^{\frac{1}{2}}dt\]

OpenStudy (anonymous):

\[ \frac{t^{\frac{3}{2} +1 }}{\frac{3}{2} +1 } - 4 \frac{t^{\frac{1}{2} +1 }}{\frac{1}{2}+1}\]

OpenStudy (anonymous):

\[\frac{2}{5} \times t^{\frac{5}{2}} - 4 \times \frac{2}{3} \times t^{\frac{3}{2}}\]

OpenStudy (anonymous):

\[\frac{2}{5} \times (4-x)^{\frac{5}{2}} - \frac{8}{3} \times (4-x)^{\frac{3}{2}}\]

OpenStudy (anonymous):

ah I am sure this is correct